# Protecting Federated Learning from Extreme Model Poisoning Attacks via Multidimensional Time Series Anomaly Detection

### v2

@@ Appendix F Datasets, Tasks, and FL Models
In Table 5, we report the full details of our experimental setup concerning the datasets used, their associated tasks, and the models (along with their hyperparameters) trained on the simulated FL environment, i.e., Flower [^11] https://flower.dev/ (see Appendix 13). Table 6 shows the description of the hyperparameters.
Table 5: Experimental setup: datasets, tasks, and FL models considered.
@@
Table 6: Description of hyperparameters.
@@
Appendix G FL Simulation Environment
@@
Table 7: Main properties of our FL environment simulated on Flower.
Total N. of Clients $K=100$
N. of Selected Clients (at each round) $m=K=100$
Ratio of Malicious Clients $r=\{0,0.2,0.6,0.8\}$
Total N. of FL Rounds $T=50$
Autoregressive Order of MAR $w=1$
Historical Window Size of Past FL Rounds $l=2$
Non-IID Dataset Distribution across Clients $\alpha_{D}=0.5$
@@
We report the main properties of our FL environment simulated on Flower in Table 7.
Appendix H Attack Settings
Below, we describe the critical parameters for each attack considered, which are also summarized in Table 8.
LIE. This method has no parameters to set.
AGR-MM. In addition to the threshold $\tau$, AGR-MM uses the so-called perturbation vectors ($\nabla^{p}$) in combination with the scaling coefficient $\gamma$ to optimize. We set $\tau=10^{-5}$ and $\nabla^{p}=\{std\}$, which is the vector obtained by computing the parameters’ inverse of the standard deviation. For Krum, we set $\nabla^{p}=\{uv\}$, which is the inverse unit vector perturbation.
Table 8: Key parameter settings for each attack strategy considered.
Attack Parameters
GAUSS $\sigma=10$
LIE N/A
OPT $\tau=10^{-5}$
AGR-MM $\tau=10^{-5};\nabla^{p}=\{uv,std\};\gamma=5$
Appendix I Defense Settings
@@
Table 9 summarizes the values of the key parameters discussed above and those characterizing our method FLANDERS.
Table 9: Key parameter settings for each defense strategy considered.
@@
Appendix J Additional Results
J.1 Malicious Detection Accuracy
Tables 10, and 11 illustrate the values for the precision ($P$) and recall ($R$) values of FLANDERS under extreme attack scenarios, integrate the findings already reported in Table 1 in the main body.
Table 10: Precision ($P$) and Recall ($R$) of FLDetector and FLANDERS in detecting malicious clients across various datasets and attacks ($r=0.6$; $T=50$ rounds).
FLDetector FLANDERS
Dataset GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
$P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$
Table 11: Precision ($P$) and Recall ($R$) of FLDetector and FLANDERS in detecting malicious clients across various datasets and attacks ($r=0.8$; $T=50$ rounds).
FLDetector FLANDERS
Dataset GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
$P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$
J.2 Aggregation Robustness Lift
First of all, Tables 12 and 13 illustrate the comparison between FLANDERS and its main competitor, FLDetector, when both are paired with standard FedAvg, under severe ($r=0.6$) and extreme ($r=0.8$) attack scenarios. These results complete those shown in Table 2 of the main submission.
In Table 14, we report the results for the CIFAR-100 dataset, which, due to space limitations, were not included in Table 3 of the main submission.
Table 12: Accuracy of the global model using FedAvg with FLDetector and FLANDERS ($r=0.6$).
Attack FedAvg FLDetector + FedAvg FLANDERS + FedAvg
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
Table 13: Accuracy of the global model using FedAvg with FLDetector and FLANDERS ($r=0.8$).
Attack FedAvg FLDetector + FedAvg FLANDERS + FedAvg
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
@@
Table 14: Accuracy of the global model on CIFAR-100 using all the baseline aggregation methods without and with FLANDERS ($r=0.8$). The best results are typed in boldface.
Aggregation CIFAR-100
GAUSS LIE OPT AGR-MM
FedAvg $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.09}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.10}$
FedMedian $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.10}$
TrimmedMean $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.11}$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$
Multi-Krum $0.08$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.11}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.11}$
Bulyan N/A N/A N/A N/A
+ FLANDERS ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.11}$
DnC $0.0.1$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.12}$
Furthermore, Tables 15, 16, 17, and 18 illustrate the impact of FLANDERS on the global model’s accuracy under light ($r=0.2$) and moderate ($r=0.6$) attack settings.
Table 15: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS on MNIST and Fashion-MNIST ($r=0.2$). The best results are typed in boldface.
Aggregation MNIST Fashion-MNIST
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.18$ $0.12$ ${\bf 0.63}$ $0.34$ $0.25$ $0.10$ $0.56$ $0.16$
+ FLANDERS ${\bf 0.86}$ ${\bf 0.83}$ $0.62$ ${\bf 0.85}$ ${\bf 0.69}$ ${\bf 0.64}$ ${\bf 0.58}$ ${\bf 0.63}$
FedMedian ${\bf 0.81}$ $0.57$ ${\bf 0.61}$ $0.63$ $0.70$ $0.63$ ${\bf 0.68}$ $0.62$
+ FLANDERS ${\bf 0.81}$ ${\bf 0.86}$ $0.58$ ${\bf 0.85}$ ${\bf 0.71}$ ${\bf 0.73}$ $0.64$ ${\bf 0.72}$
TrimmedMean $0.78$ $0.57$ ${\bf 0.76}$ $0.48$ $0.65$ $0.58$ $0.61$ $0.55$
+ FLANDERS ${\bf 0.83}$ ${\bf 0.82}$ $0.75$ ${\bf 0.77}$ ${\bf 0.69}$ ${\bf 0.69}$ ${\bf 0.63}$ ${\bf 0.70}$
Multi-Krum $0.73$ ${\bf 0.86}$ $0.83$ ${\bf 0.80}$ ${\bf 0.66}$ ${\bf 0.67}$ $0.65$ $0.64$
+ FLANDERS ${\bf 0.87}$ $0.85$ ${\bf 0.85}$ $0.77$ $0.65$ $0.64$ ${\bf 0.69}$ ${\bf 0.66}$
Bulyan $0.81$ ${\bf 0.88}$ ${\bf 0.85}$ $0.84$ ${\bf 0.67}$ ${\bf 0.72}$ ${\bf 0.70}$ ${\bf 0.76}$
+ FLANDERS ${\bf 0.83}$ ${\bf 0.88}$ $0.84$ ${\bf 0.87}$ $0.62$ $0.66$ $0.66$ $0.68$
DnC $0.20$ $0.14$ ${\bf 0.69}$ $0.30$ $0.25$ $0.18$ $0.59$ $0.17$
+ FLANDERS ${\bf 0.88}$ ${\bf 0.87}$ $0.65$ ${\bf 0.88}$ ${\bf 0.61}$ ${\bf 0.64}$ ${\bf 0.62}$ ${\bf 0.66}$
Table 16: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS on CIFAR-10 and CIFAR-100 ($r=0.2$). The best results are typed in boldface.
Aggregation CIFAR-10 CIFAR-100
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.10$ $0.10$ $0.23$ $0.10$ $0.01$ $0.01$ $0.01$ $0.02$
+ FLANDERS ${\bf 0.38}$ ${\bf 0.37}$ ${\bf 0.28}$ ${\bf 0.36}$ ${\bf 0.07}$ ${\bf 0.05}$ ${\bf 0.05}$ ${\bf 0.06}$
FedMedian ${\bf 0.34}$ $0.24$ ${\bf 0.24}$ $0.21$ $0.01$ $0.04$ $0.01$ $0.03$
+ FLANDERS ${\bf 0.34}$ ${\bf 0.28}$ $0.17$ ${\bf 0.32}$ $0.10$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.11}$
TrimmedMean ${\bf 0.35}$ $0.24$ ${\bf 0.23}$ $0.23$ $0.01$ $0.03$ $0.01$ $0.03$
+ FLANDERS ${\bf 0.35}$ ${\bf 0.33}$ $0.22$ ${\bf 0.35}$ $0.11$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.10}$
Multi-Krum $0.36$ $0.35$ $0.27$ $0.34$ $0.07$ $0.06$ $0.07$ $0.07$
+ FLANDERS ${\bf 0.42}$ ${\bf 0.44}$ ${\bf 0.35}$ ${\bf 0.42}$ ${\bf 0.10}$ ${\bf 0.09}$ ${\bf 0.10}$ ${\bf 0.09}$
Bulyan $0.40$ $0.39$ $0.34$ $0.39$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.08}$
+ FLANDERS ${\bf 0.43}$ ${\bf 0.43}$ ${\bf 0.36}$ ${\bf 0.42}$ $0.07$ $0.06$ $0.06$ ${\bf 0.08}$
DnC $0.11$ $0.10$ $0.27$ $0.15$ $0.01$ $0.01$ $0.01$ $0.02$
+ FLANDERS ${\bf 0.43}$ ${\bf 0.44}$ ${\bf 0.34}$ ${\bf 0.43}$ ${\bf 0.07}$ ${\bf 0.07}$ ${\bf 0.07}$ ${\bf 0.08}$
Table 17: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS on MNIST and Fashion-MNIST ($r=0.6$). The best results are typed in boldface.
Aggregation MNIST Fashion-MNIST
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.18$ $0.12$ $0.16$ $0.15$ $0.28$ $0.10$ $0.18$ $0.10$
+ FLANDERS ${\bf 0.78}$ ${\bf 0.82}$ ${\bf 0.79}$ ${\bf 0.85}$ ${\bf 0.69}$ ${\bf 0.65}$ ${\bf 0.65}$ ${\bf 0.62}$
FedMedian ${\bf 0.84}$ $0.18$ $0.12$ $0.16$ $0.70$ $0.10$ $0.13$ $0.10$
+ FLANDERS $0.73$ ${\bf 0.81}$ ${\bf 0.83}$ $0{\bf.82}$ ${\bf 0.72}$ ${\bf 0.73}$ ${\bf 0.71}$ ${\bf 0.72}$
TrimmedMean $0.22$ $0.12$ $0.21$ $0.13$ $0.33$ $0.10$ $0.18$ $0.10$
+ FLANDERS ${\bf 0.76}$ ${\bf 0.83}$ ${\bf 0.83}$ ${\bf 0.78}$ ${\bf 0.72}$ ${\bf 0.70}$ ${\bf 0.71}$ ${\bf 0.68}$
Multi-Krum $0.85$ $0.24$ $0.27$ $0.12$ $0.65$ $0.10$ $0.11$ $0.10$
+ FLANDERS ${\bf 0.89}$ ${\bf 0.84}$ ${\bf 0.84}$ ${\bf 0.88}$ ${\bf 0.71}$ ${\bf 0.70}$ ${\bf 0.68}$ ${\bf 0.67}$
Bulyan N/A N/A N/A N/A N/A N/A N/A N/A
+ FLANDERS ${\bf 0.86}$ ${\bf 0.87}$ ${\bf 0.81}$ ${\bf 0.88}$ ${\bf 0.65}$ ${\bf 0.69}$ ${\bf 0.66}$ ${\bf 0.67}$
DnC $0.19$ $0.10$ $0.33$ $0.24$ $0.28$ $0.10$ $0.12$ $0.10$
+ FLANDERS ${\bf 0.89}$ ${\bf 0.84}$ ${\bf 0.88}$ ${\bf 0.85}$ ${\bf 0.66}$ ${\bf 0.65}$ ${\bf 0.68}$ ${\bf 0.66}$
Table 18: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS on CIFAR-10 and CIFAR-100 ($r=0.6$). The best results are typed in boldface.
Aggregation CIFAR-10 CIFAR-100
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.10$ $0.10$ $0.10$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.36}$ ${\bf 0.36}$ ${\bf 0.34}$ ${\bf 0.35}$ ${\bf 0.08}$ ${\bf 0.09}$ ${\bf 0.08}$ ${\bf 0.09}$
FedMedian ${\bf 0.34}$ $0.10$ $0.11$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS $0.31$ ${\bf 0.31}$ ${\bf 0.30}$ ${\bf 0.33}$ $0.11$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.10}$
TrimmedMean $0.12$ $0.10$ $0.10$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.35}$ ${\bf 0.28}$ ${\bf 0.33}$ ${\bf 0.29}$ $0.10$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.12}$
Multi-Krum $0.10$ $0.35$ $0.10$ $0.13$ $0.09$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.41}$ ${\bf 0.40}$ ${\bf 0.40}$ ${\bf 0.40}$ ${\bf 0.12}$ ${\bf 0.12}$ ${\bf 0.11}$ ${\bf 0.10}$
Bulyan N/A N/A N/A N/A N/A N/A N/A N/A
+ FLANDERS ${\bf 0.43}$ ${\bf 0.41}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.09}$ ${\bf 0.08}$ ${\bf 0.10}$ ${\bf 0.08}$
DnC $0.10$ $0.10$ $0.10$ $0.11$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.09}$ ${\bf 0.09}$ ${\bf 0.09}$ ${\bf 0.10}$
Finally, in Table 19, we also assess the impact of FLANDERS in an attack-free scenario (i.e., when $r=0$). In this setting, no clear winning strategy emerges. Sometimes, FLANDERS has a detrimental effect on the global model’s accuracy with a standard aggregation mechanism (e.g., see FedAvg with the MNIST dataset). In other instances, however, FLANDERS improves the global model’s accuracy when paired with robust aggregation schemes (e.g., see Bulyan with the MNIST dataset).
Table 19: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS ($r=0$). The best results are typed in boldface.
Aggregation MNIST Fashion-MNIST CIFAR-10 CIFAR-100

### v3

@@ XII Datasets, Tasks, and FL Models
In Table XI, we report the full details of our experimental setup concerning the datasets used, their associated tasks, and the models (along with their hyperparameters) trained on the simulated FL environment, i.e., Flower [^11] https://flower.dev/ (see Appendix XIII). Table XII shows the description of the hyperparameters.
TABLE XI: Experimental setup: datasets, tasks, and FL models considered.
@@
TABLE XII: Description of hyperparameters.
@@
XIII FL Simulation Environment
@@
We report the main properties of our FL environment simulated on Flower in Table II.
XIV Defense Settings
@@
Table XIII summarizes the values of the key parameters discussed above and those characterizing our method FLANDERS.
TABLE XIII: Key parameter settings for each defense strategy considered.
@@
XV Additional Results
XV-A Impact on Attack-Free scenarios
First of all, in Table XIV, we assess the impact of FLANDERS in an attack-free scenario (i.e., when $r=0$). In this setting, no clear winning strategy emerges. Sometimes, FLANDERS has a detrimental effect on the global model’s accuracy with a standard aggregation mechanism (e.g., see FedAvg with the MNIST dataset). In other instances, however, FLANDERS improves the global model’s accuracy when paired with robust aggregation schemes (e.g., see Bulyan with the MNIST dataset).
@@ XV-A Impact on Attack-Free scenarios
TABLE XIV: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS ($r=0$). The best results are typed in boldface.
MNIST Fashion-MNIST CIFAR-10 CIFAR-100
@@
XV-B Random Client Selection
TABLE XV: Accuracy of the global model on the MNIST dataset using FLANDERS + FedAvg, with $b=2$ attackers out of $K=10$ clients, under AGR-MM attack, and with variable clients selected for training each round ($c$).
$c$ ${\bf 0.2}$ ${\bf 0.5}$ ${\bf 1.0}$
Accuracy $0.75$ $0.83$ $0.92$
The results in Table XV show that the accuracy of the global model improves as the ratio of sampled clients increases, while FLANDERS remains robust to AGR-MM attacks.
XV-C Malicious Detection Accuracy
Table XVI illustrates the values for the precision ($P$) and recall ($R$) values of FLANDERS under extreme attack scenarios, integrating the findings already reported in Tables IV, and VI in the main body.
TABLE XVI: Precision ($P$) and Recall ($R$) of FLDetector and FLANDERS in detecting malicious clients across various datasets and attacks ($r=0.6$; $T=50$ rounds).
FLDetector FLANDERS
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
$P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$
XV-D Aggregation Robustness Lift
Table XVII illustrates the comparison between FLANDERS and its main competitor, FLDetector, when both are paired with standard FedAvg, under a severe ($r=0.6$) attack scenario. These results complete those shown in Table V, and XVII of the main submission.
TABLE XVII: Accuracy of the global model using FedAvg with FLDetector and FLANDERS ($r=0.6$).
FedAvg FLDetector + FedAvg FLANDERS + FedAvg
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
Finally, Tables XVIII, and XIX illustrate the impact of FLANDERS on the global model’s accuracy under light ($r=0.2$) and severe ($r=0.6$) attack settings.
TABLE XVIII: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS ($r=0.2$). The best results are typed in boldface.
MNIST Fashion-MNIST CIFAR-10 CIFAR-100
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.18$ $0.12$ ${\bf 0.63}$ $0.34$ $0.25$ $0.10$ $0.56$ $0.16$ $0.10$ $0.10$ $0.23$ $0.10$ $0.01$ $0.01$ $0.01$ $0.02$
+ FLANDERS ${\bf 0.86}$ ${\bf 0.83}$ $0.62$ ${\bf 0.85}$ ${\bf 0.69}$ ${\bf 0.64}$ ${\bf 0.58}$ ${\bf 0.63}$ ${\bf 0.38}$ ${\bf 0.37}$ ${\bf 0.28}$ ${\bf 0.36}$ ${\bf 0.07}$ ${\bf 0.05}$ ${\bf 0.05}$ ${\bf 0.06}$
FedMedian ${\bf 0.81}$ $0.57$ ${\bf 0.61}$ $0.63$ $0.70$ $0.63$ ${\bf 0.68}$ $0.62$ ${\bf 0.34}$ $0.24$ ${\bf 0.24}$ $0.21$ $0.01$ $0.04$ $0.01$ $0.03$
+ FLANDERS ${\bf 0.81}$ ${\bf 0.86}$ $0.58$ ${\bf 0.85}$ ${\bf 0.71}$ ${\bf 0.73}$ $0.64$ ${\bf 0.72}$ ${\bf 0.34}$ ${\bf 0.28}$ $0.17$ ${\bf 0.32}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.11}$
TrimmedMean $0.78$ $0.57$ ${\bf 0.76}$ $0.48$ $0.65$ $0.58$ $0.61$ $0.55$ ${\bf 0.35}$ $0.24$ ${\bf 0.23}$ $0.23$ $0.01$ $0.03$ $0.01$ $0.03$
+ FLANDERS ${\bf 0.83}$ ${\bf 0.82}$ $0.75$ ${\bf 0.77}$ ${\bf 0.69}$ ${\bf 0.69}$ ${\bf 0.63}$ ${\bf 0.70}$ ${\bf 0.35}$ ${\bf 0.33}$ $0.22$ ${\bf 0.35}$ ${\bf 0.11}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.10}$
Multi-Krum $0.73$ ${\bf 0.86}$ $0.83$ ${\bf 0.80}$ ${\bf 0.66}$ ${\bf 0.67}$ $0.65$ $0.64$ $0.36$ $0.35$ $0.27$ $0.34$ $0.07$ $0.06$ $0.07$ $0.07$
+ FLANDERS ${\bf 0.87}$ $0.85$ ${\bf 0.85}$ $0.77$ $0.65$ $0.64$ ${\bf 0.69}$ ${\bf 0.66}$ ${\bf 0.42}$ ${\bf 0.44}$ ${\bf 0.35}$ ${\bf 0.42}$ ${\bf 0.10}$ ${\bf 0.09}$ ${\bf 0.10}$ ${\bf 0.09}$
Bulyan $0.81$ ${\bf 0.88}$ ${\bf 0.85}$ $0.84$ ${\bf 0.67}$ ${\bf 0.72}$ ${\bf 0.70}$ ${\bf 0.76}$ $0.40$ $0.39$ $0.34$ $0.39$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.08}$
+ FLANDERS ${\bf 0.83}$ ${\bf 0.88}$ $0.84$ ${\bf 0.87}$ $0.62$ $0.66$ $0.66$ $0.68$ ${\bf 0.43}$ ${\bf 0.43}$ ${\bf 0.36}$ ${\bf 0.42}$ $0.07$ $0.06$ $0.06$ ${\bf 0.08}$
DnC $0.20$ $0.14$ ${\bf 0.69}$ $0.30$ $0.25$ $0.18$ $0.59$ $0.17$ $0.11$ $0.10$ $0.27$ $0.15$ $0.01$ $0.01$ $0.01$ $0.02$
+ FLANDERS ${\bf 0.88}$ ${\bf 0.87}$ $0.65$ ${\bf 0.88}$ ${\bf 0.61}$ ${\bf 0.64}$ ${\bf 0.62}$ ${\bf 0.66}$ ${\bf 0.43}$ ${\bf 0.44}$ ${\bf 0.34}$ ${\bf 0.43}$ ${\bf 0.07}$ ${\bf 0.07}$ ${\bf 0.07}$ ${\bf 0.08}$
TABLE XIX: Accuracy of the global model using all the baseline aggregation methods without and with FLANDERS ($r=0.6$). The best results are typed in boldface.
MNIST Fashion-MNIST CIFAR-10 CIFAR-100
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.18$ $0.12$ $0.16$ $0.15$ $0.28$ $0.10$ $0.18$ $0.10$ $0.10$ $0.10$ $0.10$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.78}$ ${\bf 0.82}$ ${\bf 0.79}$ ${\bf 0.85}$ ${\bf 0.69}$ ${\bf 0.65}$ ${\bf 0.65}$ ${\bf 0.62}$ ${\bf 0.36}$ ${\bf 0.36}$ ${\bf 0.34}$ ${\bf 0.35}$ ${\bf 0.08}$ ${\bf 0.09}$ ${\bf 0.08}$ ${\bf 0.09}$
FedMedian ${\bf 0.84}$ $0.18$ $0.12$ $0.16$ $0.70$ $0.10$ $0.13$ $0.10$ ${\bf 0.34}$ $0.10$ $0.11$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS $0.73$ ${\bf 0.81}$ ${\bf 0.83}$ $0{\bf.82}$ ${\bf 0.72}$ ${\bf 0.73}$ ${\bf 0.71}$ ${\bf 0.72}$ $0.31$ ${\bf 0.31}$ ${\bf 0.30}$ ${\bf 0.33}$ ${\bf 0.11}$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.10}$
TrimmedMean $0.22$ $0.12$ $0.21$ $0.13$ $0.33$ $0.10$ $0.18$ $0.10$ $0.12$ $0.10$ $0.10$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.76}$ ${\bf 0.83}$ ${\bf 0.83}$ ${\bf 0.78}$ ${\bf 0.72}$ ${\bf 0.70}$ ${\bf 0.71}$ ${\bf 0.68}$ ${\bf 0.35}$ ${\bf 0.28}$ ${\bf 0.33}$ ${\bf 0.29}$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.12}$
Multi-Krum $0.85$ $0.24$ $0.27$ $0.12$ $0.65$ $0.10$ $0.11$ $0.10$ $0.10$ $0.35$ $0.10$ $0.13$ $0.09$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.89}$ ${\bf 0.84}$ ${\bf 0.84}$ ${\bf 0.88}$ ${\bf 0.71}$ ${\bf 0.70}$ ${\bf 0.68}$ ${\bf 0.67}$ ${\bf 0.41}$ ${\bf 0.40}$ ${\bf 0.40}$ ${\bf 0.40}$ ${\bf 0.12}$ ${\bf 0.12}$ ${\bf 0.11}$ ${\bf 0.10}$
Bulyan N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
+ FLANDERS ${\bf 0.86}$ ${\bf 0.87}$ ${\bf 0.81}$ ${\bf 0.88}$ ${\bf 0.65}$ ${\bf 0.69}$ ${\bf 0.66}$ ${\bf 0.67}$ ${\bf 0.43}$ ${\bf 0.41}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.09}$ ${\bf 0.08}$ ${\bf 0.10}$ ${\bf 0.08}$
DnC $0.19$ $0.10$ $0.33$ $0.24$ $0.28$ $0.10$ $0.12$ $0.10$ $0.10$ $0.10$ $0.10$ $0.11$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.89}$ ${\bf 0.84}$ ${\bf 0.88}$ ${\bf 0.85}$ ${\bf 0.66}$ ${\bf 0.65}$ ${\bf 0.68}$ ${\bf 0.66}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.42}$ ${\bf 0.09}$ ${\bf 0.09}$ ${\bf 0.09}$ ${\bf 0.10}$